The units digit is 1 more than 3 times the tens digit in a two digit number. The number represented when the digits are interchanged is 8 times the sum of the digits. Find the original number.
Answer by ankor@dixie-net.com(22740)

The units digit is 1 more than 3 times the tens digit in a two digit number.
The number represented when the digits are interchanged is 8 times the sum of the digits.
Find the original number.
:
Let x = the 10's digit
Let y = the units
then
10x+y = the number
:
Write an equation for each statement:
:
"The units digit is 1 more than 3 times the tens digit"
y = 3x + 1
:
"The number represented when the digits are interchanged is 8 times the sum of the digits."
10y + x = 8(x+y)
:
You should be able to solve it now